OpenStudy (anonymous):

A copy machine makes 60 copies per minute. a second machine makes 80 copies per minute. the second machine starts making copies 2 minutes after the first machine starts. both machines stop making copies 8 minutes after the first machine started. together, the 2 machines made how many copies?

OpenStudy (anonymous):

\[\frac{ 60 copies }{ \min }*(8\min) +\frac{ 80copies }{ \min }*(8-2)\min=Total\]\[480 copies + 480 copies= 960 copies\]